# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

### v1

@@
Abstract.
The observations documented in Cyber Threat Intelligence (CTI) reports play a critical role in describing adversarial behaviors, providing valuable insights for security practitioners to respond to evolving threats. Recent advancements of Large Language Models (LLMs) have demonstrated significant potential in various cybersecurity applications, including CTI report understanding and attack knowledge graph construction. While previous works have proposed benchmarks that focus on the CTI extraction ability of LLMs, the sequential characteristic of adversarial behaviors within CTI reports remains largely unexplored, which holds considerable significance in developing a comprehensive understanding of how adversaries operate. To address this gap, we introduce AttackSeqBench, a benchmark tailored to systematically evaluate LLMs’ capability to understand and reason attack sequences in CTI reports. Our benchmark encompasses three distinct Question Answering (QA) tasks, each task focuses on the varying granularity in adversarial behavior. To alleviate the laborious effort of QA construction, we carefully design an automated dataset construction pipeline to create scalable and well-formulated QA datasets based on real-world CTI reports. To ensure the quality of our dataset, we adopt a hybrid approach of combining human evaluation and systematic evaluation metrics. We conduct extensive experiments and analysis with both fast-thinking and slow-thinking LLMs, while highlighting their strengths and limitations in analyzing the sequential patterns in cyber attacks. The overarching goal of this work is to provide a benchmark that advances LLM-driven CTI report understanding and fosters its application in real-world cybersecurity operations. Our dataset and code are available at https://github.com/Javiery3889/AttackSeqBench.
Cyber Security, Cyber Threat Intelligence, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
1. Introduction
Figure 1. Illustration of the difference between existing CRU-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020). We highlight the sequential patterns within CTI reports and investigate the capability of LLMs in comprehending and reasoning such sequential patterns.
The increasing complexity and diversity of cyber attacks have imposed a challenge to defend against Advanced Persistent Threats (APTs) (Huang et al., 2018; Duo et al., 2022). Cyber Threat Intelligence (CTI) reports, which detail the adversarial behaviors observed in cyber attacks, are crucial towards addressing emerging cyber threats. However, their unstructured nature demands extensive manual analysis (Wagner et al., 2019; Sun et al., 2023). Meanwhile, recent advancements in Large Language Models (LLMs) have demonstrated promising applications in cybersecurity, such as attack knowledge graph construction (Zhang et al., 2025b; Hu et al., 2024) and threat intelligence extraction (Rahman et al., 2023; Xu et al., 2024b). These capabilities shed new light on LLM-driven CTI Report Understanding (CRU) to support cybersecurity operations, leading to the development of various CRU-related benchmarks to assess LLMs’ capability in the field.
Existing CRU-related benchmarks primarily focus on evaluating LLMs on three aspects: (1) Classifying adversarial behaviors (Alam et al., 2023) (2) Extracting threat intelligence from unstructured text (Bhusal et al., 2024), or (3) Understanding meaningful correlations between the cyber threats and malware groups (Alam et al., 2024). While these benchmarks focus on the information extraction capabilities of LLMs, their ability to understand and reason about the sequential patterns of adversarial behavior remains largely unexplored. We refer to these sequences as attack sequences (Choi et al., 2021; Al-Sada et al., 2025), which represent the temporal order of malicious actions adversaries perform to achieve their goals. The actions which constitute an attack sequence are commonly referred to as attack patterns, characterizing adversarial behavior on a tactical level (Mavroeidis and Bromander, 2021; Siracusano et al., 2023). We illustrate the key differences between existing CRU-related tasks and analyzing attack sequences in Figure 1. Although knowledge bases (KBs) such as the MITRE ATT&CK® framework (Strom et al., 2018) document real-world adversary behaviors through defined attack patterns, analyzing the patterns individually is insufficient to fully capture the progression of cyber attacks as documented in CTI reports. The sophisticated and stealthy nature of APTs requires a comprehensive understanding of how adversaries transition between the different attack phases, which are orchestrated as an attack sequence. This raises the need to consider the sequential characteristics of a cyber attack during CRU.
To address this research gap, we delve into the suitability of LLMs in analyzing attack sequences. We highlight three key perspectives when formulating our main research task:
Perspective 1: The process of mining threat intelligence from CTI reports not only requires general text comprehension; it also necessitates expert-level cybersecurity knowledge. This requirement becomes apparent in attack sequence analysis, as it requires a deep understanding of adversarial behavior to formulate the associate attack sequence. Although LLMs demonstrate significant potential in CRU, prior research has identified limitations in their factual reliability, particularly in technical and knowledge-intensive tasks (Wang et al., 2023a; Xu et al., 2024a). In real-world cybersecurity operations, these limitations can hinder precise and timely analysis of CTI reports, ultimately impacting defensive strategies.
Perspective 2: LLMs have demonstrated remarkable capabilities across a broad range of scenarios, from traditional Natural Language Processing (NLP) tasks to more even more challenging real-world problems (Zhao et al., 2023). The recent breakthrough of slow-thinking reasoning LLMs has also shed new light on complex reasoning tasks related to coding and mathematical reasoning (DeepSeek-AI, 2025; OpenAI, 2025). Yet, the extent to which LLMs can be applied to attack sequence analysis has remained under-explored.
Perspective 3: To address the constantly evolving cyber threat landscape and the advancements of LLMs, the benchmark must be extensible to incorporate newly observed attack sequences from CTI reports. Additionally, the unstructured nature of CTI reports necessitates the benchmark to be generalizable to ensure a comprehensive and realistic evaluation of LLMs’ capabilities. Most importantly, the benchmark must be reproducible to facilitate fair and consistent evaluations across the different LLMs.
Building on these perspectives, we introduce AttackSeqBench, a novel benchmark designed for extensive investigation. Specifically, considering Perspective 1, we define three distinct Question Answering (QA) benchmarking tasks that follow the three-level structure of adversary behaviors —Tactics, Techniques, and Procedures (TTPs) —as outlined in the ATT&CK KB. This hierarchical representation allows security practitioners to represent observed attack patterns in varying levels of granularity based on pre-defined adversarial behaviors. Each tactic represents a high-level objective within a cyber attack’s lifecycle, while techniques and procedures elaborate on the specific actions performed by an adversary to achieve their objectives. Following Perspective 2, we design three benchmark settings to extensively evaluate the effectiveness and robustness of AttackSeqBench. Moreover, we leverage a diverse set of fast-thinking and slow-thinking LLMs in our experiments (Min et al., 2024), providing comprehensive insights into the capabilities and limitations of LLMs in attack sequence analysis. Catering to Perspective 3, we construct attack sequences based on real-world CTI reports. Additionally, we carefully design an automated dataset generation pipeline to generate questions in our QA dataset based on the attack sequence constructed. We summarize our contributions as follow:
  • We propose AttackSeqBench, a novel benchmark designed to systematically evaluate LLMs’ capabilities in analyzing attack sequences constructed from real-world CTI reports.
  • We introduce three QA tasks in our benchmark, each aligned with the different levels of granularity in adversarial behavior representation.
  • We conduct extensive experiments and various benchmarking settings using a diverse set of fast-thinking and slow-thinking LLMs to provide insights into their capabilities and limitations.
2. Related Work
CTI Report Understanding and LLMs. CTI refers to the collection and analysis of evidence-based information related to cyber attacks and threat actors, enabling security practitioners to gain actionable insights into the adversarial operations respond to these adversaries effectively (Wagner et al., 2019; Sun et al., 2023). A crucial source of CTI comes from CTI reports published by security vendors, these reports are typically presented as unstructured textual documents and document observations based on emerging and ongoing cyber threats (Rahman et al., 2020). CTI Report Understanding (CRU) broadly encompasses the extraction, interpretation, and reasoning of key insights within CTI reports. TTP extraction is a critical aspect in CRU, enabling security practitioners to characterize adversarial behaviors into known attack patterns defined in the ATT&CK KB (Husari et al., 2017). Some works (Legoy et al., 2020; Marchiori et al., 2023) leverage multi-label text classification to extract TTPs, while more recent research utilize Language Models (LMs) to perform TTP extraction (Rani et al., 2024). Another line of work (Shin et al., 2023) explores embedding TTPs using Global Vectors for Word Representation to capture co-occurrences. Despite considerable progress, these works tend to be costly and error-prone as adversarial behavior continues evolving.
With the breakthrough of LLMs, researchers have progressively explored their applicability within the cybersecurity domain (Zhang et al., 2024a). Within CRU, existing research (Fayyazi et al., 2024) has showcased significant capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data. Another line of works (Huang and Xiao, 2024; Cheng et al., 2024) utilizes LLMs to extract the relevant entities to construct knowledge graphs (KGs) based on the CTI reports. AttacKG+ (Zhang et al., 2025b) constructs adversarial behavior KGs by leveraging LLMs to map extracted entities to ATT&CK TTPs. However, the capabilities of LLMs to analyze attack sequences in CTI reports remain unexplored.
Related Benchmarks. Within the cybersecurity domain, existing work (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023) have proposed several benchmarks to evaluate LLMs’ cybersecurity knowledge. However, these works primarily rely on authoritative sources (e.g., textbooks, open standards, and example TTPs from ATT&CK KB) and do not incorporate CTI reports in their dataset. CTIBench (Alam et al., 2024) designs four CTI-related tasks, including threat actor attribution and vulnerability assessment, yet only the threat actor attribution task leverages CTI reports. CTISum (Peng et al., 2024) evaluates the ability of LLMs to summarize key facts and attack processes from CTI reports. Extending this work, SevenLLM (Ji et al., 2024) designs an additional task that focuses on extracting entities and relations. To add on, prior research (Siracusano et al., 2023) assesses LLMs’ capabilities in extracting attack patterns, along with entities and relations. In the general domain, existing works (Wang et al., 2023b; Valmeekam et al., 2023) have proposed benchmarks to evaluate the reasoning capabilities of LLMs in sequence-related tasks. Furthermore, previous research has highlighted LLMs capabilities in temporal event analysis (Zhang et al., 2024b; Chang et al., 2024). However, these works may not accurately reflect the extent to which LLMs can reason about attack sequences. In contrast, AttackSeqBench delves into the suitability of LLMs in attack sequence analysis through three tasks that requires in-depth knowledge of adversarial behavior, while emphasizing the practical aspects of CRU by utilizing real-world CTI reports to construct attack sequences in our dataset.
3. Task Definition and Dataset
3.1. Task Definition
An attack sequence $S$ models the progression of adversarial behavior in a cyber attack, characterized by the logical order of the ATT&CK tactics observed in a CTI report alongside the associated techniques and procedures. Formally, $S$ is defined as a 4-tuple:
$S=(T,E,P,O),$
where:
  1. (1)
    Tactic Sequence $T=(t_{1},\ldots,t_{n})$: An ordered list of ATT&CK tactics, which outlines the phases of a cyber attack identified in a CTI report. Note that tactic $t_{k}$ represents the $k$-th tactic.
  2. (2)
    Technique Mapping $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$: A set of ATT&CK techniques (and/or sub-techniques) associated with tactic $t_{k}$. The overall set of techniques across all tactics in the CTI report is represented as $E=\bigcup_{k=1}^{n}E(t_{k})$.
  3. (3)
[…]

### v2

# AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence

@@
Abstract
Cyber Threat Intelligence (CTI) reports document observations of cyber threats, synthesizing evidence about adversaries’ actions and intent into actionable knowledge that informs detection, response, and defense planning. However, the unstructured and verbose nature of CTI reports poses significant challenges for security practitioners to manually extract and analyze such sequences. Although large language models (LLMs) exhibit promise in cybersecurity tasks such as entity extraction and knowledge graph construction, their understanding and reasoning capabilities towards behavioral sequences remains underexplored. To address this, we introduce AttackSeqBench, a benchmark designed to systematically evaluate LLMs’ reasoning abilities across the tactical, technical, and procedural dimensions of adversarial behaviors, while satisfying Extensibility, Reasoning Scalability, and Domain-dpecific Epistemic Expandability. We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across the proposed 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain. Our findings contribute to a deeper understanding of LLM-driven CTI report understanding and foster its application in cybersecurity operations.
1 Introduction
Amid rapid digital transformation, the increasing sophistication and diversity of cyber attacks have become a pervasive concern for cybersecurity globally (Duo et al., 2022). Cyber Threat Intelligence (CTI) reports, which document observations of these threats, have emerged as a crucial resource in proactive defenses (Wagner et al., 2019). However, they are often lengthy and unstructured, resulting in a labor-intensive task for practitioners to manually analyze and extract insights (Sun et al., 2023).
Recently, Large Language Models (LLMs) have demonstrated promising potential in several cybersecurity applications (Zhang et al., 2024a). This sheds new light towards incorporating LLMs into CTI Report Understanding (CRU) task, where we define CRU as a broad concept encompassing tasks that derive and reason threat intelligence from CTI reports. However, existing benchmarks primarily assess LLMs on threat intelligence extraction and attack attribution, while their potential for understanding adversarial behaviors dependencies in CTI reports remains largely unexplored (cf. Appendix A.7). Such ability is crucial in anticipating future malicious attack actions, particularly in multi-stage cyber attacks launched by Advanced Persistent Threats (APTs) (Li et al., 2022).
As illustrated in Figure 1, we define the sequence of adversary behaviors as attack sequence (Al-Sada et al., 2025) to represent the execution flow of malicious actions across different stages of a cyber attack under the MITRE ATT&CK® framework (Strom et al., 2018). Building on this definition and the following key perspectives, we further delve into the suitability of LLMs in analyzing attack sequences. 1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate attack sequences from newly observed CTI reports. 2) Reasoning Scalability: Recently, Large Reasoning Models (LRMs) have demonstrated substantial advantages over conventional LLMs in multi-step reasoning tasks, such as coding and mathematical reasoning. However, existing CRU works have primarily focused on addressing CTI–related tasks via LLM, leaving the necessity of reasoning for attack sequence analysis largely unexplored. 3) Domain-Specific Epistemic Expandability: LLMs exhibited limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024b), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations. This requirement becomes particularly pronounced in attack sequence analysis, which necessitates a comprehensive understanding of adversarial behaviors to effectively reason multi-stage cyber attacks.
Figure 1: Illustration an example cyber attack sequence and our AttackSeqBench.
Building upon these perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis. Catering to Extensibility, we first construct attack sequences based on extensive real-world CTI reports, ensuring that the benchmark accurately reflects the complexity and diversity of Tactics, Techniques, and Procedures (TTPs) in cyber attacks performed by APTs. Moreover, we design three Question Answering (Q&A) tasks under the adversary behaviors hierarchy in MITRE ATT&CK® and develop an automated Q&A generation pipeline that converts newly-collected CTI reports into the pre-defined format, enabling its extensibility on the corpus side. Following Reasoning Scalability, we further evaluate several LRMs and reasoning distillation strategies, which function well in general domains, to identify their strengths and limitations on the specialized attack sequence analysis task, providing helpful insights for future research in this area. To achieve Domain-Specific Epistemic Expandability, we aggregate cybersecurity-related knowledge from some existing benchmarks and embed it into LLMs via several post-training strategies to examine their epistemic expandability at the model level. Moreover, we also extend beyond the conventional zero-shot setting by introducing context-based and RAG-empowered settings, which pertinently assess LLMs’ epistemic expandability when injecting domain-specific cybersecurity knowledge at the semantic and representation levels.
Our contribution are as follows: (I) We introduce AttackSeqBench, the pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to analyze attack sequences across diverse inference settings and multi-level tasks. (cf. Section 2) (II) We quantitatively demonstrate that existing LRMs fail to substantially outperform LLMs on attack sequence analysis and perform markedly worse in most cases, a contrast to their advantages observed in domains such as mathematics and coding. (cf. Section 3.3) (III) We offer a comprehensive analysis of how parameterization and parameter scale affect existing models’ attack sequence analysis, and further examine why current LRMs and RAG underperform on this specialized task. This work uncovers the fundamental limitations of current models in attack sequence analysis and provides actionable insights to guide future research in this domain. (cf. Section 3.4 and Section 3.5)
2 Dataset Construction and Verification
Goal of our AttackSeqBench Our goal is to explore the capability of diverse types of LLMs in attack sequence understanding. Through comprehensive evaluation across various tasks and settings, we emphasize the strengths and limitations of existing LLMs, offering the promising yet underexplored directions.
2.1 Problem Definitions
CTI report understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge. To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB. Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
    Tactic Sequence $T$: An ordered list of ATT&CK tactics, such that $T=(t_{1},\ldots,t_{n})$, where $t_{k}$ is the $k$-th tactic in the sequence.
    Technique Mappings $E$: The set of ATT&CK techniques / sub-techniques in $S$, where $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$ denote all the techniques / sub-techniques that belong to tactic $t_{k}$.
[…]
    CTI Outline $O$: A textual summary of the organized TTPs based on the order of Tactic Sequence $T$, such that $O=(o_{1},\ldots,o_{n})$, where $o_{k}$ refers to the summarized text associated with tactic $t_{k}$.
Figure 2: Overview of our automated QA dataset construction pipeline.
2.2 Dataset Construction
As illustrated in Figure 2, we first construct attack sequences using the extracted TTPs and CTI outline from CTI reports. Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the QA dataset.
Attack Sequence Construction. To efficiently and massively extract threat intelligence from these unstructured reports, we utilize a set of 408 CTI reports from various security vendors (Cisco Talos Intelligence Group, 2025; Microsoft, 2025) to construct attack sequences that accurately reflect the behaviors of real-world APTs. Specifically, we utilize a LLM-based KG construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combin them to construct the attack sequences $S$. Notably, we exclude CTI outlines which contains less than two ATT&CK tactics in attack sequence construction as they are unlikely to detail attack patterns observed in real-world cyber attacks.
Q&A Generation. Inspired by the remarkable question generation abilities of LLMs across multiple domains (Alam et al., 2024; Zhang et al., 2024b; Mucciaccia et al., 2025), we adopt an answer-aware question generation approach using GPT-4o (OpenAI, 2024a). To elaborate, we first instruct the LLM to generate a seed Q&A pair for each tactic, technique, and group of procedures with the given attack sequence. Furthermore, we utilize the model’s In-Context Learning ability to generate the more relevant Q&A pairs (Dong et al., 2024), by including the CTI outline and few-shot Q&A examples in the question generation prompt (cf. Appendix C.1).
For the Multiple-Choice Question (MCQ) tasks, we adopt a rule-based approach to select three choices as distractors. Specifically, we select a adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k+1}$ or $t_{k-1}$) and randomly select two tactics from the ATT&CK KB in AttackSeq-Tactic. Regarding AttackSeq-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first technique belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\notin E(t_{k})$; (2) The second technique is supported by the given attack sequence but belongs to another tactic $e_{j,j}\notin E(t_{k})$; (3) The third technique comes from a randomly chosen tactic that is not supported by the given attack sequence.
Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeq-Procedure-Yes. Next, we randomly sample its 70% questions to generate the negative question samples. Specifically, we design two types of Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” and/or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
Q&A Refinement. While LLMs possess remarkable text generation capabilities, these models may deviate from the requirements specified in users’ instructions (Joshi et al., 2025), resulting in the conflict between the generated questions and the order of TTPs in attack sequences. Inspired by the Self-Refine framework (Madaan et al., 2023), we design a refinement criteria to iteratively refine the initial questions via the same LLM. To perform a holistic evaluation, we introduce six aspects below that emphasizes the question’s linguistic (i.e., Clarity) and task-oriented properties (Fu et al., 2024). Here, we divide the task-oriented aspects into three categories: (1) Question Complexity (i.e., Answerability); (2) Content Alignment (i.e., Relevance, Consistency, Answer Consistency); (3) Attack Sequence Alignment (i.e., Logical) (cf. Appendix A.2).
Considering the foundational role of Answerability in benchmark design, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives. Questions failing this requirement are discarded from the next step. Secondly, the LLM is instructed to evaluate the questions based on the remaining five aspects, providing a numerical score (out of five) and feedback for each aspect. Lastly, the LLM is prompted to refine the questions based on the feedback given (cf. Appendix C.2). We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final QA dataset.
After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/1,393/3,249 to 1,697/1,917/1,223/1,412, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria. Additionally, we further illustrate the top-10 ATT&CK tactics and techniques within our dataset in Figure 7(a) and 7(b) respectively. The most frequent tactic and technique in the figure is associated with a key objective of APTs, highlighting the relevance of our Q&A dataset in capturing attack sequences based on real-world cyber attacks.
Table 1: Evaluation results of four sub-tasks on human evaluation and automatic evaluation.
Task Human Evaluation Automatic Evaluation
Hum. Perf. Ans. Cla. Log Rel. Con. Ans. Con. Ans. Cla. Log Rel. Con. Ans. Con.
Tactic 0.51 4.30 4.36 4.45 4.56 4.46 4.44 4.52 4.65 4.79 4.84 4.65 4.76
Technique 0.71 4.09 4.21 4.40 4.45 4.44 4.41 4.10 4.40 4.62 4.63 4.39 4.59
Procedure-Yes 0.74 4.88 4.70 4.88 5.00 4.81 4.94 4.02 4.06 4.61 4.47 3.78 3.89
Procedure-No 0.56 4.55 4.84 - - 4.82 4.66 3.29 3.66 - - 2.77 3.25
Average 0.63 4.45 4.53 4.57 4.67 4.63 4.61 3.98 4.19 4.67 4.64 3.90 4.12
2.3 Dataset Evaluation
LLMs demonstrate strong potential in solving complex tasks, but they inevitably exhibit even severer hallucinations, which has become a widely recognized concern in the research community. To address this, we adopt a hybrid approach towards evaluating the quality of the constructed QA dataset using the criteria defined in our Q&A refinement (cf. Section 2.2). We design 5-point Likert scales for each of the evaluation criterion, where higher scores indicate better alignment.
Human Evaluation. We first randomly sample 35 questions from each sub-task to construct a question set for human evaluation. Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 2.2. Based on Table 1, we observe that the average Human Performance (abbreviated as Hum. Perf.) equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise. Notably, AttackSeq-Procedure-No is derived from AttackSeq-Procedure-Yes through Yes-to-No transferring strategies, that is, its Logical and Relevance are inherently misaligned with the attack sequence, and we therefore do not evaluate these two aspects. Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.45$ to $4.61$ out of $5$, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
Automatic Evaluation. To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain, including cybersecurity (Xu et al., 2024a). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated Q&A pairs. Specifically, we design individual prompts for each aspect in the evaluation criteria that includes its definition and the scoring guideline based on the same 5-point Likert scale in Human Evaluation (cf. Table LABEL:tab.annotation_instructions in Appendix A.2). Then we instruct GPT-4o rate the Q&A for each aspect based on the evaluation criterion and the correct answer from the ATT&CK KB (cf. Appendix C.3). Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences (cf. ). The fact that automatic evaluation scores are lower than human evaluation scores further indicates that answering questions correctly in our AttackSeqBench is more challenging for LLMs than for domain experts.
3 Benchmark and Experiments
Figure 3: Overview of the three benchmark settings that exhibit varying levels of contextual information given to the LLM.
3.1 Benchmark Settings
As illustrated in Figure 3, we elaborate on three benchmark settings with varying levels of contextual information: (1) Zero-shot setting, (2) Context setting, and (3) RAG-empowered setting.
Zero-shot Setting. Motivated by the significant zero-shot reasoning abilities of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and the Q&A pairs to evaluate the LLMs’ performance on three tasks based on their inherent knowledge.
Context Setting. Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize this context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level. Here, we remove the corresponding summary of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$. Afterwards, the LLM will be instructed to use the corresponding $O_{m}$ to answer the question, highlighting its potential to perform abductive reasoning to determine the most plausible TTP in attack sequence.
RAG-empowered Setting. Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability pf LLMs and mitigate hallucinations (Zhang et al., 2025a). Here, we also design the RAG-empowered setting to evaluate the Domain-Specific Epistemic Scalability of LLMs at the representation level. This leverages the LLMs’ in-context learning ability to learn the associations between the entity in the question’s body and the relevant TTPs, thereby decomposing the problem and eliciting its stronger reasoning ability (Wu et al., 2022).
3.2 Implementation Details
To investigate the CRU capability of existing models, we evaluate seven LLMs (i.e., LLaMa-3.1-8B (Grattafiori et al., 2024), ChatGLM-4-9B (GLM et al., 2024), Qwen-2.5-3B, Qwen-2.5-14B, Qwen-2.5-32B (QwenTeam, 2024), Llama-3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (Llama-3.1-8B), R1(Qwen-2.5-14B), R1(Qwen-2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench. We also utilize four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed security knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of our AttackSeqBench (cf. Appendix A.4). Here, we measure the performance with accuracy Acc=n/MAcc\!=\!n/M, where $n$ is the correctly-answered number of questions and $M$ is the total number.  [^1] Due to the page limitation, we introduce the complete implementation details in Appendix A.5 and A.6.
Table 2: Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
Zero-Shot Context RAG Zero-Shot Context RAG Zero-Shot Context RAG
[…]
3.3 Performance Comparison
Comparison between diverse groups of LLMs. As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of the LLMs consistently outperforms the others, and the optimal LLM varies diverse tasks. For instance, the best-performing models under the zero-shot setting across the three benchmark tasks are Qwen-2.5-32B, Qwen-2.5-32B, and GPT-o3-mini, respectively. This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pretraining corpus is likely overshadowed by general-domain content. Moreover, most models consistently perform worst in AttackSeq-Tactic compared to the other two tasks, mirroring the human evaluation results in Section 2.3 and underscoring the common challenge faced by both human experts and LLMs in tactical inference.
Furthermore, from Table 3, we can observe that compared to the zero-shot setting, all models exhibit substantial performance gains on AttackSeq-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences. As defined in Appendix A.3, AttackSeq-Procedure-No is inherently more complex and reasoning-demanding than AttackSeq-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence. This explains why LRMs with stronger reasoning ability outperform in AttackSeq-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability. Finally, most post-training strategies substantially improve the performance of its base LLM, particularly in zero-shot settings that rely solely on internal knowledge. However, their performance still lags behind instructive LLMs equipped with task-adapted prompts. This highlights a promising direction: Designing specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity.
Table 3: Performance comparison of two LLMs and two LRMs on AttackSeq-Procedure-Yes and AttackSeq-Procedure-No in each benchmark setting, where the bold value indicates the best performance of each column.
LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
Zero-Shot Regular RAG Zero-Shot Regular RAG
Llama-3.1-8B 0.9128 0.7572 0.8858 0.2434 0.6216 0.2111
GPT-4o 0.9469 0.9567 0.8831 0.4426 0.2698 0.4143
R1 (Llama-8B) 0.9332 0.8427 0.9191 0.1508 0.4417 0.1792
GPT-o3-mini 0.7612 0.7048 0.7408 0.6303 0.6678 0.5552
Comparison on Contextual Information. Comparing the performance of LLMs across three benchmark settings in Table 2, we can observe that: In general, the Context setting consistently outperforms Zero-Shot and RAG settings across most benchmark tasks, with the advantage more pronounced in larger LLMs. Taking the Qwen-2.5 series as an example: performance shifts from zero-shot being optimal in Qwen-2.5-3B (0.4467 vs. 0.4614 vs. 0.3296) to context-setting being optimal in Qwen-2.5-32B (0.6195 vs. 0.5903 vs. 0.5154) in the AttackSeq-Tactic task, with Qwen-2.5-14B showing the transition in between. This phenomenon is reasonable as larger LLMs possess more extensive internal knowledge, and task-specific context further enhances their effectiveness and robustness within the specific domain. Moreover, both LLMs and LRMs consistently fail to reach optimal performance under the RAG-empowered setting. This indicates that naive retrieval integration may introduce additional noise instead of enhancing results, underscoring the requirement for more advanced retrieval-augmented approaches. We further investigate its limitation in Section 3.5.2.
3.4 Robustness Analysis
Figure 4: Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeq-Tactic under the zero-shot setting.
3.4.1 Parameter Sensitivity Analysis
Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and illustrate them in Figure 4(a) and Figure 4(b) respectively. Firstly, we observe that increasing the temperature from 0 to 1 causes a sharp performance drop in smaller LLMs, while larger LLMs remain relatively unaffected in attack sequence analysis. This may be because smaller LLMs lack discriminative power and oscillate among suboptimal answers, whereas larger ones generate more stable logits that preserve correct outputs even under smoothing, while aligning with the observations of previous work in general domain (Renze, 2024; Li et al., 2025). On the other hand, we also releaze that increasing the token budget yields stable performance and output length for LLMs, whereas LRMs achieve significant gains in both performance and output tokens. Specifically, when the Max Output Tokens increase from 1,024 to 4,096, R1 (LLaMA-8B) and R1 (Qwen-32B) improve accuracy by 13.29% and 16.74%, with average output tokens increasing by 37.35% and 43.28%, respectively. However, when the token budget is further increased to 8,192, LRMs exhibit diminishing returns: average output tokens increase by 13.27% and 9.83%, while accuracy improves only by 0.29% and 0.46%. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in LRMs while considering the associated costs incurred (Wang et al., 2024).
3.4.2 Computational Complexity Analysis
Figure 5: Computational complexity analysis of seven LLMs and four LRMs in AttackSeq-Tactic under the regular setting. The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
[…]